Respected sir,
[1] 94 is divided in to two parts in such a way that the fifth part of the first and the eighth part of the second are in the ratio 3:4,the first part is ?
[2] Rs.1980 are divided among A,B,C so that half of A's part ,one-third of B's part and one-sixth of C's part are equal.Then , B's part is ?
plz solve my ratio ,proportion difficualty .waiting for yur reply.
thank you sir.
Regards.
Binita.

Found 2 solutions by solver91311, edjones:
Answer by solver91311(12126)


Let x represent the first part of 94. Then 94 - x is the second part. The fifth part of the first part is: , and the eighth part of the second part is: .

These two numbers are in the ratio 3:4, so:



Cross-multiply:





Cross-multiply:







Check: The fifth part of 30 is 6, and the eighth part of 64 is 8 and 6 is to 8 as 3 is to 4.

********************************************

Half of A's part is equal to one-third of B's part:



Half of A's part is equal to one-sixth of C's part:



Then:













Check:

Half of 360 is 180, one-third of 540 is 180, and one-sixth of 1080 is 360. 360 + 540 + 1080 = 1980.

John


Answer by edjones(7311)

Let f=first part and s=the second
f+s=94
f=94-s
.
.2f/.125s=3/4
.2(94-s)/.125s=3/4
(18.8-.2s)/.125s=3/4
4(18.8-.2s)=3*.125s
75.2-.8s=.375s
1.175s=75.2
s=64
f=30
I'll try to get back to you regarding the 2nd part.
.
Ed